Solve.
The U.S population in 1990 was approximately 250 million, and the average growth rate for the past 30 years gives a doubling time of 66 years. The above formula for the United States then becomes (p(in millions) = 250 X 2 to the power of (y-1990)/66
What will the population of the United States be in 2025 if this growth rate continues?

Answer by stanbon(75887)

(p(in millions) = 250 X 2 to the power of (y-1990)/66
What will the population of the United States be in 2025 if this growth rate continues?
-----------
P(2025) = 250*2^[(2025-1990)/66] in millions
P(2025) = 250*2^[0.530303...] in millions
P(2025) = 250*1.444232517.. in millions
P(2025) = 361.058129... in millions
=======================
Cheers,
Stan H.